OpenStudy (anonymous):

Please help soon! I really don't understand!Jake and Amanda drove for 9 hours against a wind speed of 8 mph. Traveling with the same wind speed, it took them 7 hours to make the return trip. How fast were they traveling in still air? 64 mph 66 mph 68 mph 70 mph

OpenStudy (anonymous):

\[\Delta x = V \Delta t\] delta x is the same for the trip there and back. [distance] \[\Delta x = (V - 8)*9\] \[\Delta x = (V + 8)*7\] so \[(V+8)*7 = (V - 8)*9\] 7V + 56 = 9V - 72 Solve for V

OpenStudy (anonymous):

So you use that formula to get the answer?

OpenStudy (anonymous):

I am getting all different answers for everything.

OpenStudy (anonymous):

its a mix of the formula on the first line + logic which is specific per problem. I get V = 64

OpenStudy (anonymous):

So 64? Thank you so much!

OpenStudy (anonymous):

glad i could help

OpenStudy (anonymous):

:)